Simon Company’s year-end balance sheets follow.

At December 31 Current Yr 1 Yr Ago 2 Yrs Ago
Assets
Cash $ 29,254 $ 34,195 $ 35,264
Accounts receivable, net 89,400 62,900 51,400
Merchandise inventory 112,000 83,000 59,000
Prepaid expenses 9,421 8,976 3,918
Plant assets, net 255,755 238,369 206,618
Total assets $ 495,830 $ 427,440 $ 356,200
Liabilities and Equity
Accounts payable $ 123,462 $ 72,237 $ 47,018
Long-term notes payable secured by
mortgages on plant assets
93,216 98,311 78,720
Common stock, $10 par value 162,500 162,500 162,500
Retained earnings 116,652 94,392 67,962
Total liabilities and equity $ 495,830 $ 427,440 $ 356,200

The company’s income statements for the Current Year and 1 Year Ago, follow. Assume that all sales are on credit:

For Year Ended December 31 Current Yr 1 Yr Ago
Sales $ 644,579 $ 508,654
Cost of goods sold $ 393,193 $ 330,625
Other operating expenses 199,819 128,689
Interest expense 10,958 11,699
Income tax expense 8,380 7,630
Total costs and expenses 612,350 478,643
Net income $ 32,229 $ 30,011
Earnings per share $ 1.98 $ 1.85


Compute days' sales uncollected.

Days' Sales Uncollected
Choose Numerator: / Choose Denominator: x Days = Days' Sales Uncollected
/ x = Days' Sales Uncollected
Current Yr: / x = 0 days
1 Yr Ago: / x = 0 days

Compute accounts receivable turnover.

Accounts Receivable Turnover
Choose Numerator: / Choose Denominator: = Accounts Receivable Turnover
/ = Accounts receivable turnover
Current Yr: / = times
1 Yr Ago: / = times
Answer #1

Accounts receivable turnover = net credit sales / Average accounts receivable

Current year = 644579 /(89400+62900)/2 = 8.46 times

1 year ago = 508654 /(62900+51400)/2 = 8.69 days

Days sales uncollected = 365/ accounts receivable turnover

Current year = 365 / 8.46 = 43.12 days

1 year ago = 365 /8.69 = 42 days